how do i solve this quadratic eqaution 7x+36=x+4
Found 2 solutions by rfer, richard1234:
Answer by rfer(16322)

This is not a quadratic
6x=-32
x=5.33

Answer by richard1234(7193)

That is not a quadratic, since it only has degree 1. It can still be solved for x using algebraic methods:

7x + 36 = x + 4

6x = -32

x = -32/6 = -16/3.